OpenStudy (4meisu):

Find the first term of the sequence 5,10,20,40, ... which exceeds 10.000

OpenStudy (4meisu):

*10,000

OpenStudy (anonymous):

\(10\times 2^k=10,000\) solve for \(k\)

OpenStudy (anonymous):

The expression appears to be: \[(2.5)2^n\]So we look for an \(n\) such that: \[ (2.5)2^n>10000 \]Which is \[ 2^n>\frac{10000}{2.5}\implies \log_2(2^n)=n>\log_2(\frac{10000}{2.5}) \]